# Analyzing the factors that are involved in length of inpatient stay at the hospital for diabetes patients

### v1

@@ 1. Introduction
In recent years, the surge in diabetes cases has raised significant concerns, intensifying the demand for medical resources, (wang2021trends, ) and (fang2021trends, ). This is a crucial endeavor amidst the mounting challenges posed by the increasing diabetic population and the lingering effects of the COVID-19 pandemic. Current research aims to construct a model that quantifies the factors influencing inpatient hospital stay durations for diabetes patients, enabling hospital administrators to gain a deeper understanding of patient needs and enhance their management strategies.
@@ 2. Literature Review
Predictive modeling plays a crucial role in managing diabetes-related hospitalizations. (alturki2019predictors, ) addressed the growing concern of unplanned hospital readmissions and extended hospital stays, suggesting that early prediction of these outcomes could facilitate targeted interventions and improved resource allocation. (koziol2021predictors, ) investigated risk factors associated with rehospitalization and mortality following diabetes-related hospital admissions in Poland, further illustrating the need for robust predictive models. Scholars have utilized various machine learning techniques to predict hospital stays. (barsasella2022machine, ) demonstrated that linear regression works well in Length of Stay (LoS) prediction. Advanced techniques, such as the stacked ensemble method discussed by (alahmar2018application, ), highlight the potential for data-driven strategies to improve healthcare management and outcomes for diabetes patients. Recent advancements, such as the work by (xu2024kernel, ), on time series analysis provide deeper insights into the temporal aspects of patient data.
Over the past three years, several COVID-19 outbreaks have caused varying degrees of bed shortages, especially ICU beds, in both developed countries with ample medical resources and developing countries with already poor medical conditions (sen2021closer, ). Sometimes, the COVID-19 testing process itself, rather than the infection, delays patient discharge due to requirements for negative test results before patients can be transferred to nursing care facilities (olanipekun2021impact, ). This often results in an additional 2-3 days of hospital stay before discharge. The pandemic has also underscored the importance of advanced modeling techniques. (xu2022multi, ) discussed the application of multi-view kernel clustering to explore the nonlinear relationships in healthcare data, which is crucial for understanding the complex interactions between COVID-19 and diabetes management.
Recent advancements in kernel-based methods offer promising tools for quantifying factors affecting hospital stay durations for diabetes patients by uncovering nonlinear relationships (xu2024rhine, ; pethunachiyar2020classification, ; xu2018self, ). The latest model, Kolmogorov-Arnold Networks (KAN), introduces a deep learning framework with high interpretability, offering a promising strategy for analyzing complex, time-dependent healthcare data (xu2024kolmogorov, ). This concept of interpretability is similar to that of (nwegbu2022novel, ; xu2024rhine, ), who emphasized the importance of dynamic regime discovery in forecasting states of data (e.g., healthcare data), and offers new ways to improve the predictive accuracy of diabetes management.
@@ 3.4.1. Exploratory Data Analysis and Feature Engineering
The factor variable levels were reordered so that the most frequent level was first. After the changes 9,997 records remained in the dataset.
To explore the data, here exhibits some descriptive statistics.(AppendixA)

### v2

@@ 1. Introduction
In recent years, the surge in diabetes cases has raised significant concerns, intensifying the demand for medical resources, (wang2021trends, ) and (fang2021trends, ; xu2023drnet, ). This is a crucial endeavor amidst the mounting challenges posed by the increasing diabetic population and the lingering effects of the COVID-19 pandemic. Current research aims to construct a model that quantifies the factors influencing inpatient hospital stay durations for diabetes patients, enabling hospital administrators to gain a deeper understanding of patient needs and enhance their management strategies.
@@ 2. Literature Review
Predictive modeling plays a crucial role in managing diabetes-related hospitalizations. (alturki2019predictors, ) addressed the growing concern of unplanned hospital readmissions and extended hospital stays, suggesting that early prediction of these outcomes could facilitate targeted interventions and improved resource allocation. (koziol2021predictors, ) investigated risk factors associated with rehospitalization and mortality following diabetes-related hospital admissions in Poland, further illustrating the need for robust predictive models. Scholars have utilized various machine learning techniques to predict hospital stays. (barsasella2022machine, ) demonstrated that linear regression works well in Length of Stay (LoS) prediction. Advanced techniques, such as the stacked ensemble method discussed by (alahmar2018application, ), highlight the potential for data-driven strategies to improve healthcare management and outcomes for diabetes patients. Recent advancements, such as the work by (xu2024kernel, ; xu2025towards, ), on time series analysis provide deeper insights into the temporal aspects of patient data.
Over the past three years, several COVID-19 outbreaks have caused varying degrees of bed shortages, especially ICU beds, in both developed countries with ample medical resources and developing countries with already poor medical conditions (sen2021closer, ; xu2025drift2matrix, ). Sometimes, the COVID-19 testing process itself, rather than the infection, delays patient discharge due to requirements for negative test results before patients can be transferred to nursing care facilities (olanipekun2021impact, ). This often results in an additional 2-3 days of hospital stay before discharge. The pandemic has also underscored the importance of advanced modeling techniques. (xu2022multi, ; xu2020kernel, ; chen2021self, ) discussed the application of multi-view kernel clustering to explore the nonlinear relationships in healthcare data, which is crucial for understanding the complex interactions between COVID-19 and diabetes management.
Recent advancements in kernel-based methods offer promising tools for quantifying factors affecting hospital stay durations for diabetes patients by uncovering nonlinear relationships (xu2024rhine, ; pethunachiyar2020classification, ; xu2018self, ). The latest model, Kolmogorov-Arnold Networks (KAN), introduces a deep learning framework with high interpretability, offering a promising strategy for analyzing complex, time-dependent healthcare data (xu2024kolmogorov, ; chen2022clustering, ; xu2024kan4drift, ; chen2022dynamic, ; xu2024kan, ). This concept of interpretability is similar to that of (nwegbu2022novel, ; xu2024rhine, ; xu2024wormhole, ), who emphasized the importance of dynamic regime discovery in forecasting states of data (e.g., healthcare data), and offers new ways to improve the predictive accuracy of diabetes management.
@@ 3.4.1. Exploratory Data Analysis and Feature Engineering
The factor variable levels were reordered so that the most frequent level was first. After the changes 9,997 records remained in the dataset. To explore the data, here exhibits some descriptive statistics (AppendixA).